OpenStudy (anonymous):

PERCENT OF CHANGE? 360 pounds to 352 pounds

OpenStudy (anonymous):

8/360 = 0.022 = 2.22%

jimthompson5910 (jim_thompson5910):

Percent Change = [(New - Old)/Old] * 100 Percent Change = [(352-360)/360] * 100 Percent Change = [-8/360] * 100 Percent Change = -0.02222 * 100 Percent Change = -2.222% Since we have a negative change, there is a percent decrease of 2.222%